OpenStudy (anonymous):

What is the distance between points P(-3, 5) and Q(3, 1)?

OpenStudy (arindameducationusc):

use distance formula

OpenStudy (arindameducationusc):

52

OpenStudy (arindameducationusc):

just wait.. I will show you how....

OpenStudy (arindameducationusc):

See this attachment

OpenStudy (arindameducationusc):

https://www.ltcconline.net/greenl/courses/154/factor/circle.htm

OpenStudy (arindameducationusc):

this link is for more info

OpenStudy (anonymous):

\[\sqrt{(3-(3))^{2}+(1-5)^{2}}= 2\sqrt{13} = 7.211\]

OpenStudy (anonymous):

thanks for you help! ill have to practice more

OpenStudy (anonymous):

Looks like I'm missing a negative sign in the (3-(-3)) part

OpenStudy (arindameducationusc):

yes @saseal . Guys is my answer correct?

OpenStudy (anonymous):

You forgot to square root the 52, otherwise you are good.

OpenStudy (anonymous):

\[\sqrt{52} = 2\sqrt{13}\]

OpenStudy (arindameducationusc):

O.... I forgot the underroot...... @saseal Yes,yes.....